User name Sparveris
Log entry time 00:01:20 on May25,2007
Entry number 203139
16:00 Shift takes over during run 3736 (currently running Q2=68
on 3He @ 90uA).
16:20 MCC called - beam will be off for some time to do some optics. Run 3736 ended (350K).
18:42 Beam is back @ 80uA. Starting new run 3737.
Beam is not alignet at both BPMs - asked them to set it to both
at (-2.1,1) - they are doing it through run 3737 (its a short
run, we will keep on taking some data in the meantime.
run 3737 = Test run.
19:14 Got acceptable & aligned beam @ 70uA. Prescales were not
set properly during the morning shift (data taking from 13:00-16:00,
runs 3730-3736). Prescales are set ok now. Started good production
run under stable beam conditions (run 3738).
22:25 Changed PS3 to 2 (from 1). Started run 3742. deadtime went
from 7% to 1%.
22:41 Lost Q2/Right - ended run 3742. Got Tech-on-call, trying to figure out the problem.
23:45 Q2 Reset - asked for beam again. Started run 3743.
00:00 New shift takes over during run 3743.